# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ 5 Conclusions
Looking ahead, dynamic scoring approaches could be explored to capture the evolving nature of XAI techniques and their use in real-world applications. In conclusion, PASTA intends to take a step towards creating a transparent and trustworthy AI ecosystem. By aligning AI explanations with human cognitive processes, we aim to foster the development of more interpretable AI systems that can be understood and trusted by users across various domains.
Appendix A PASTA-dataset: process
A.1 Classifier training
A.1.1 Dataset
@@
The dataset used in PASTA is designed to provide a benchmark for evaluating a wide range of XAI techniques across different explanation modalities. To ensure robustness and versatility, we have built a benchmark dataset consisting of four diverse, publicly available datasets, each bringing distinct characteristics in terms of visual content and concept annotations. Choosing which task to focus on is a tough question. We have chosen to focus on the image classification task. This task can be performed in many different domains, but in order not to be too domain-specific, we decided to work on general datasets. These datasets enable the evaluation of both image-based and concept-based XAI methods.
The dataset used to train our inference model integrates four datasets, each chosen for its unique attributes that align with the requirements of evaluating explainability methods. The datasets are as follows:
@@
  • COCO: A widely-used dataset known for its complexity and variety, containing 117k training images, 4.5k validation images annotated with 80 object categories, which we consider to be concepts in the images. The task we focus on is indoor scene labeling, to do so, we took the subset of images of indoor scenes (53,051 images). Then, we labeled the images using a scene label DNN trained on the MIT SUN.
@@
  • Cats Dogs Cars: A curated dataset featuring images of cats, dogs, and cars. The goal of this dataset is to explore if color biases are present in the model or not. It is composed of 3,858 training images, 39 concepts, and 3 classes. Since this network does not include annotated concepts, we used Grounding DINO (Liu et al., 2023) as an annotator.
@@
  • Image-level annotations: These are traditional class labels or object categories that describe the primary content of the image.
@@
This dual-level annotation setup ensures that XAI methods can be evaluated not only for their ability to explain class predictions but also for how well they handle concept-based explanations. The presence of both granular (part-level, concept) and holistic (object, scene-level) annotations provides a comprehensive evaluation environment for various XAI methods.
@@
A.1.2 Specific procedures for CBMs and black boxes models.
To explain the various training procedures for our CBMs, we decompose them into two components: the concept extractor and the classifier. The concept extractor generates an embedding from an input image, with each element representing a concept, while the classifier predicts the label from this embedding. We categorize the CBMs we use based on the training methods for these two components. For CLIP-based CBMs (LaBo, CLIP-linear, and CLIP-QDA), the concept extraction is performed in a zero-shot manner i.e., we only use the training images and labels to train the classifier. For CBMs that require training the concept extractor, we use the concept annotations provided by each dataset.
@@ A.1.3 Results
A.2 XAI techniques
Table 6 presents an overview of the different XAI methods integrated into the dataset. A brief description of each method is provided below to summarize their key features and mechanisms.
@@ A.2 XAI techniques
BCos (Böhle et al., 2024) Interpretable latent space Image Ante-hoc ResNet50-BCos
@@
GradCAM (Selvaraju et al., 2017) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
HiResCAM (Draelos & Carin, 2020) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
GradCAMElementWise (Pillai & Pirsiavash, 2021) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
GradCAM++ (Chattopadhay et al., 2018) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
XGradCAM (Fu et al., 2020) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
AblationCAM (Ramaswamy et al., 2020) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
ScoreCAM (Wang et al., 2020) Perturbation Image Post-hoc ViT, ResNet50
@@
EigenCAM (Muhammad & Yeasin, 2020) Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
EigenGradCAM (Muhammad & Yeasin, 2020) Gradient+Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
LayerCAM (Jiang et al., 2021) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
FullGrad (Srinivas & Fleuret, 2019) Gradient Image Post-hoc ViT, ResNet50
@@
Deep Feature Factorizations (Collins et al., 2018) Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
SHAP (Lundberg & Lee, 2017) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
LIME (Ribeiro et al., 2016) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
X-NeSyL (Díaz-Rodríguez et al., 2022) Interpretable latent space Concepts Ante-hoc X-NeSyL
@@
CLIP-linear-sample (Yan et al., 2023) Interpretable latent space Concepts Ante-hoc CLIP-linear
CLIP-QDA-sample (Kazmierczak et al., 2024) Counterfactual Concepts Ante-hoc CLIP-QDA
LIME-CBM (Kazmierczak et al., 2024) Perturbation Concepts Post-hoc CLIP-QDA, ConceptBottleneck
SHAP-CBM (Kazmierczak et al., 2024) Perturbation Concepts Post-hoc CLIP-QDA, ConceptBottleneck
RISE (Petsiuk et al., 2018) Perturbation Concepts Post-hoc ConceptBottleneck
@@
LIME (Local Interpretable Model-agnostic Explanations): LIME explains individual predictions of any classifier by approximating it locally with an interpretable model. It perturbs the input and observes how the predictions change, identifying the most influential parts of the input for the prediction. This is a saliency-based XAI method, visualizing the important regions in an image that the DNN relies on.
@@
SHAP (SHapley Additive exPlanations): SHAP is a unified approach to interpreting model predictions based on Shapley values from cooperative game theory. It assigns each feature an importance value for a particular prediction, offering a sound measure of feature importance. This is a saliency-based XAI method, visualizing the important regions in an image that the DNN relies on.
@@
GradCAM (Gradient-weighted Class Activation Mapping): GradCAM visualizes the regions in an image that contribute to the classification. It uses the gradients of the target concept (e.g., a specific class) flowing into the final convolutional layer to produce a coarse localization map highlighting important regions. This is a saliency-based XAI method.
@@
AblationCAM: AblationCAM improves GradCAM by iteratively removing parts of the input and observing the output effect to identify important regions. This is a saliency-based XAI method, visualizing the crucial regions in an image that the DNN relies on.
@@
EigenCAM: EigenCAM applies PCA to the activations of the last convolutional layer to produce a saliency map. It highlights the directions in which activations show the most variance, identifying critical features. This is a saliency-based XAI method.
@@
FullGrad: FullGrad computes gradients of the output with respect to both the input and intermediate layer outputs, aggregating these gradients to generate a comprehensive saliency map. It is a saliency-based XAI method that visualizes key regions in an image.
@@
GradCAMPlusPlus: GradCAMPlusPlus improves GradCAM with a refined weighting scheme for the gradients, allowing better handling of multiple occurrences of the target concept. This is a saliency-based XAI method.
@@
GradCAMElementWise: GradCAMElementWise extends GradCAM by considering element-wise multiplications of gradients and activations, producing more precise visual explanations. This is a saliency-based XAI method.
@@
HiResCAM: HiResCAM improves on class activation mapping by using higher-resolution feature maps for more detailed visual explanations. This is a saliency-based XAI method.
@@
ScoreCAM: ScoreCAM improves CAM methods by using output scores to weight the activation maps’ importance, providing a more faithful saliency map without relying on gradients. This is a saliency-based XAI method.
@@
XGradCAM: XGradCAM integrates cross-layer information to combine saliency maps from different layers, producing a more comprehensive explanation. This is a saliency-based XAI method.
@@
DeepFeatureFactorization: This method decomposes feature representations learned by a deep model into interpretable factors. It provides insights into how features contribute to the model’s decisions, being a saliency-based XAI method.
@@
CLIP-QDA-sample: This model uses the CLIP framework and applies Quadratic Discriminant Analysis (QDA) for classification. It links visual and concept-based representations to provide interpretable explanations. This is a concept bottleneck model (CBM).
@@
CLIP-Linear-sample: Similar to CLIP-QDA, this model leverages the CLIP framework but applies logistic regression for classification, providing interpretable explanations based on concept representations. This is a CBM.
@@
LIME CBM: This model generates a list of concepts and applies logistic regression. It uses LIME to highlight the most important concepts for classification.
SHAP CBM: This model generates a list of concepts and applies logistic regression, using SHAP to emphasize the most crucial concepts in classification.
Labo: Labo extracts human-interpretable concepts and maps them to the model’s internal representations for more comprehensible decision-making explanations.
RISE (Randomized Input Sampling for Explanation): RISE generates heatmaps by perturbing input regions and measuring their impact on model outputs. This technique identifies the most influential regions in the model’s decision-making process.
@@
Finally, Figure 7 shows the distribution of XAI techniques applied across the datasets. To enhance the generalizability of our results, we increased the diversity of XAI techniques used. This was achieved by not applying every technique to every image uniformly, allowing for a more diverse set of explanations to be generated. This variability ensures that our analysis captures a broad spectrum of interpretability techniques, providing deeper insights into the performance of XAI techniques across different datasets and models.
@@
A.3 Dataset Annotation process
The annotation process took place via an online web application, created and deployed by a contracting company. 15 participants were recruited to take part in the annotation process. These participants ranged in age from 19 to 37 (mean age 25.9, standard deviation 5.5). Figure 8 shows the age distribution. Among the participants, 5 identified themselves as male, 10 as female, 0 as non-binary, 0 did not wish to say. All participants were based in India.

### v2

@@ Conclusions
Dynamic scoring approaches could be explored to capture the evolving nature of XAI techniques and their use in real-world applications. PASTA intends to take a step towards creating a transparent and trustworthy AI ecosystem. By aligning AI explanations with human preferences, we aim to foster the development of more interpretable AI systems that can be understood and trusted by users. This work also introduces a perceptual metric, paving the way for future research to implement the PASTA-score as a perceptual loss aimed at enhancing the trustworthiness of networks, drawing for example, inspiration from the emerging use of LPIPS (Zhang et al. 2018) in tasks such as image generation (Jo, Yang, and Kim 2020).
Acknoledgements
This work was performed using HPC resources from GENCI-IDRIS (Grant 2024 - AD011014675R1).
Appendix A A. PASTA-dataset: process
A.1 Generation of explanations
Classifier training
@@ A.1 Generation of explanations
The PASTA-dataset is designed to provide a benchmark for evaluating a wide range of XAI techniques across different explanation modalities. To ensure robustness and versatility, our dataset is based on four, publicly available datasets, each bringing distinct characteristics in terms of visual content and concept annotations. Choosing which task to focus on is a tough question. We have chosen to focus on image classification. This task can be performed in many different domains, but in order not to be too domain-specific, we decided to work on general datasets. These datasets enable the evaluation of both image-based and concept-based XAI methods.
Then, the initial phase in constructing the PASTA-dataset involves training the various classifier models on which explanations will be generated. Specifically, we utilize ResNet50 (He et al. 2016), ViT-B (Dosovitskiy 2020), ResNet50-BCos (Böhle et al. 2024), CLIP-Linear (Yan et al. 2023), CLIP-QDA (Kazmierczak et al. 2024), X-NeSyL (D´ıaz-Rodr´ıguez et al. 2022), and CBM (Koh et al. 2020). These models are trained separately on each classifier’s dataset referenced below.
It is important to note that certain classifiers utilized in our platform necessitate concept-level annotations. Specifically, for each sample, information regarding the presence or absence of each concept in the image, along with their respective bounding boxes, is required. This detailed information is not natively available in each dataset. Consequently, we enhanced each dataset to meet this requirement. The datasets and the corresponding modifications are as follows:
@@
  • COCO: A widely-used dataset known for its complexity and variety, containing 117k training images, 4.5k validation images annotated with 80 object categories, which we consider to be concepts in the images. The labels correspond for this specific dataset to indoor scene labeling, to do so, we took the subset of images of indoor scenes (53,051 images). Then, we labeled the images using a scene label DNN trained on the MIT SUN.
@@
  • Cats Dogs Cars: A curated dataset featuring images of cats, dogs, and cars. The goal of this dataset is to explore if color biases are present in the model or not. It is composed of 3,858 training images, 39 concepts, and 3 classes. Since this network does not include annotated concepts, we used Grounding DINO (Liu et al. 2023) as an annotator. Since the number of images that constitute Cate Dogs Cars is sufficiently small, we manually checked the bounding boxes generated and found no significant errors.
@@
  • Image-level annotations: These are traditional class labels (Table LABEL:table:table_classes) or object categories that describe the primary content of the image.
@@
Table 4: List of classes used in all the datasets used to train our inference models.
Dataset Labels
catsdogscars cat, dog, car
pascalpart aeroplane, bicycle, bird, bottle, bus, car, cat, cow, dog, horse, motorbike, person, pottedplant, sheep, train, tvmonitor
monumai Baroque, Gothic, Hispanic-Muslim, Renaissance
coco shopping_and_dining, workplace, home_or_hotel, transportation, sports_and_leisure, cultural
@@
As previously indicated, additional care must be taken when training CBMs. To explain the various training procedures for our CBMs, we decompose them into two components: the concept extractor and the classifier. The concept extractor generates an embedding from an input image, with each element representing a concept, while the classifier predicts the label from this embedding. We categorize the CBMs we use based on the training methods for these two components. For CLIP-based CBMs (LaBo, CLIP-linear, and CLIP-QDA), the concept extraction is performed in a zero-shot manner i.e., we only use the training images and labels to train the classifier. For CBMs that require training the concept extractor (X-NeSyL and ConceptBottleneck), we use the concept annotations provided by each dataset.
@@
Computation of explanations
Upon completion of the training process for all classifiers, the subsequent phase involves the generation of explanations for their respective inferences. A comprehensive enumeration of the XAI methods, along with the corresponding classifiers to which each method is applied, is presented in Table 5. For this, we use for the computation of explanations a subset of 250 images of the classifier’s dataset test split per dataset, that serve as the basis of the PASTA-dataset. This diverse selection ensures a broader generalization of the XAI techniques across datasets being assessed. Note that, unlike traditional datasets, our benchmark dataset comprises a triplet of images, explanations, and labels. This triplet enables us to quantitatively assess the quality of XAI techniques.
Precisely, we tested 14 saliency-based XAI techniques and 6 concept-based XAI techniques, with a particular emphasis on a variety of functioning of methods. A beneficial aspect of our approach is that many of the methods are tested on multiple backbone architectures. This aspect is particularly important, as many of the post-hoc methods evaluated in our study are designed to be DNN-agnostic. In such cases, the XAI method is applied to independently trained models. For instance, GradCAM is evaluated on ResNet50, ViT-B, and CLIP-zero-shot models, while SHAP-CBM is applied to both CLIP-QDA and the original Concept Bottleneck model as proposed by Koh et al. (2020).
@@
BCos (Böhle et al. 2024) Interpretable latent space Image Ante-hoc ResNet50-BCos
@@
GradCAM (Selvaraju et al. 2017) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
HiResCAM (Draelos and Carin 2020) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
GradCAMElementWise (Pillai and Pirsiavash 2021) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
GradCAM++ (Chattopadhay et al. 2018) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
XGradCAM (Fu et al. 2020) Gradient Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
AblationCAM (Ramaswamy et al. 2020) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
ScoreCAM (Wang et al. 2020) Perturbation Image Post-hoc ViT, ResNet50
@@
EigenCAM (Muhammad and Yeasin 2020) Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
EigenGradCAM (Muhammad and Yeasin 2020) Gradient+Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
FullGrad (Srinivas and Fleuret 2019) Gradient Image Post-hoc ViT, ResNet50
@@
Deep Feature Factorizations (Collins, Achanta, and Susstrunk 2018) Factorization Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
SHAP (Lundberg and Lee 2017) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
LIME (Ribeiro, Singh, and Guestrin 2016) Perturbation Image Post-hoc ViT, ResNet50, CLIP (zero-shot)
@@
X-NeSyL (D´ıaz-Rodr´ıguez et al. 2022) Interpretable latent space Concepts Ante-hoc X-NeSyL
@@
CLIP-linear-sample (Yan et al. 2023) Interpretable latent space Concepts Ante-hoc CLIP-linear
CLIP-QDA-sample (Kazmierczak et al. 2024) Counterfactual Concepts Ante-hoc CLIP-QDA
LIME-CBM (Kazmierczak et al. 2024) Perturbation Concepts Post-hoc CLIP-QDA, ConceptBottleneck
SHAP-CBM (Kazmierczak et al. 2024) Perturbation Concepts Post-hoc CLIP-QDA, ConceptBottleneck
RISE-CBM (Petsiuk, Das, and Saenko 2018) Perturbation Concepts Post-hoc ConceptBottleneck
A brief description of each method is provided below to summarize their key features and mechanisms.
@@
LIME (Local Interpretable Model-agnostic Explanations): LIME explains individual predictions of any classifier by approximating it locally with an interpretable model. It perturbs the input and observes how the predictions change, identifying the most influential parts of the input for the prediction.
@@
SHAP (SHapley Additive exPlanations): SHAP is a unified approach to interpreting model predictions based on Shapley values from cooperative game theory. It assigns each feature an importance value for a particular prediction, offering a sound measure of feature importance.
@@
GradCAM (Gradient-weighted Class Activation Mapping): GradCAM visualizes the regions in an image that contribute to the classification. It uses the gradients of the target concept (e.g., a specific class) flowing into the final convolutional layer to produce a coarse localization map highlighting important regions.
@@
AblationCAM: AblationCAM improves GradCAM by iteratively removing parts of the input and observing the output effect to identify important regions.
@@
EigenCAM: EigenCAM applies PCA to the activations of the last convolutional layer to produce a saliency map. It highlights the directions in which activations show the most variance, identifying critical features.
@@
FullGrad: FullGrad computes gradients of the output with respect to both the input and intermediate layer outputs, aggregating these gradients to generate a comprehensive saliency map.
@@
GradCAMPlusPlus: GradCAMPlusPlus improves GradCAM with a refined weighting scheme for the gradients, allowing better handling of multiple occurrences of the target concept.
@@
GradCAMElementWise: GradCAMElementWise extends GradCAM by considering element-wise multiplications of gradients and activations, producing more precise visual explanations.
@@
HiResCAM: HiResCAM improves on class activation mapping by using higher-resolution feature maps for more detailed visual explanations.
@@
ScoreCAM: ScoreCAM improves CAM methods by using output scores to weight the activation maps’ importance, providing a more faithful saliency map without relying on gradients.
@@
XGradCAM: XGradCAM integrates cross-layer information to combine saliency maps from different layers.
@@
DeepFeatureFactorization: This method decomposes feature representations learned by a deep model into interpretable factors. It provides insights into how features contribute to the model’s decisions.
@@
CLIP-QDA-sample: This model uses the CLIP framework and applies Quadratic Discriminant Analysis (QDA) for classification. This methodology employs counterfactuals on the conceptual representations of images to generate explanations.
@@
CLIP-Linear-sample: This model also uses the CLIP framework but employs logistic regression for classification, thereby offering interpretable explanations grounded in the transparency of the regression analysis.
@@
LIME CBM: This model generates a list of concepts and applies logistic regression. The methodology employs LIME to identify and highlight the most significant concepts at the conceptual level for classification purposes.
SHAP CBM: This model generates a list of concepts and applies logistic regression, using SHAP on the concept level to emphasize the most crucial concepts in classification.
Labo: Similar to CLIP-Linear-sample, Labo extracts human-interpretable concepts and maps them to the model’s internal representations to facilitate more comprehensible decision-making explanations. This process leverages the transparency of its classification mechanism, albeit utilizing a custom transparent network.
RISE: RISE (Randomized Input Sampling for Explanation) generates heatmaps by perturbing input regions and measuring their impact on model outputs. This technique identifies the most influential regions in the model’s decision-making process.
@@
The PASTA dataset comprises 21,100 instances, each containing images, predictions, and explanations. These instances were subsequently evaluated by human annotators, resulting in 633,000 unique Likert ratings. This extensive evaluation was achieved by asking six questions to five annotators for each instance. We aggregate these evaluations using majority voting to favor consensus opinions. This dataset size is fairly standard in the Automated Essay Scoring literature (Lee et al. 2024), where the objective is to train a model to predict human-assigned scores. Figure 12 shows the distribution of XAI techniques applied across the datasets. To enhance the generalizability of our results, we increased the diversity of XAI techniques used. This was achieved by not applying every technique to every image uniformly, allowing for a more diverse set of explanations to be generated. This variability ensures that our analysis captures a broad spectrum of interpretability techniques, providing deeper insights into the performance of XAI techniques across different datasets and models.
@@
A.2 Human evaluation protocol
Once the explanations are performed, we can quantify the interpretability and usefulness of XAI techniques accurately, using a human evaluation of the quality of explanations. Our human-centric approach complements existing approaches that focus primarily on assessing the model’s internal behavior. For example, traditional evaluations of faithfulness measure how closely an explanation corresponds to the model’s true functioning, while we assess in our dataset how the explanation fits human expectations.
Desiratas
First, we establish a comprehensive set of assessment criteria that are evaluated on a graded scale. We consolidate different criteria from the literature into the following set of desiderata for XAI explanations that we wish to evaluate:
  • Trustworthiness(Arrieta et al. 2020) measures the extent to which an explanation accurately reflects whether a model will act as intended when facing a given problem.
  • Robustness (Doshi-Velez and Kim 2017; Agarwal et al. 2022a; Yeh et al. 2019) assesses the stability and relevance of the explanation across a broad range of models and inputs.
  • Complexity (Nauta et al. 2023; Nguyen and Mart´ınez 2020; Bhatt, Weller, and Moura 2021) checks whether the explanation is both simple and informative, balancing clarity and detail.
  • Objectivity (Bennetot et al. 2022) evaluates whether the explanation is interpreted consistently by the majority within a given audience.
Evaluation Protocol
Then, we apply an evaluation protocol, developed with the help of a psychologist, to ensure that annotators fully understand the task and the expectations. This includes annotator training and close monitoring throughout the process. Interfaces, as well as the formulation of questions, play a key role in the quality of the annotations (Pommeranz et al. 2012), and their design must be considered cautiously to avoid confounding cognitive biases. The formulation of the questions has been carefully chosen to ensure that they are fully understood by each annotator. To maintain consistency and reliability, all annotators undergo a training session before starting the actual annotation task. This training familiarizes them with the XAI techniques, evaluation criteria, rating scale and datasets, ensuring a uniform understanding of the task and the expectations.
The annotation process took place via an online web application, created and deployed by a contracting company. 24 participants were recruited to take part in the annotation process. These participants ranged in age from 19 to 37 (mean age 24.58, standard deviation 3.19). Figure 13 shows the age distribution. Among the participants, 7 identified themselves as male, 17 as female, 0 as non-binary, 0 did not wish to say. All participants were based in India. Each participant’s task was to annotate 147 explanations. To address the subjectivity inherent in the task, each instance, comprising a triplet of image, explanation, and label, is evaluated by five different annotators.